OpenStudy (2jessicae2):

Write the sum using summation notation, assuming the suggested pattern continues. -9 - 3 + 3 + 9 + ... + 81

OpenStudy (lord_box):

What does it look like is happening to each term?

OpenStudy (2jessicae2):

they are adding 6

OpenStudy (lord_box):

Alright, and it starts from -9, how would you represent that using summation notation?

OpenStudy (2jessicae2):

im not sure

OpenStudy (lord_box):

\[\sum_{n=0}^{?} -9 + ?\] This is what you know so far. What could you add to the -9 to show that you're adding 6 , then 12, then 18, etc every time?

OpenStudy (2jessicae2):

6n

OpenStudy (lord_box):

That's right. Now how many times would you have to add 6n to get to 81?

OpenStudy (2jessicae2):

15

OpenStudy (lord_box):

Alright, \[\sum_{n=0}^{15} -9 + 6n\] This is how you build it in summation notation, or alternatively, \[\sum_{n=11}^{16} -9 + 6(n-1)\], same thing

OpenStudy (lord_box):

The second one should read n=1, sorry

OpenStudy (2jessicae2):

thank you so much for the help

OpenStudy (lord_box):

You're welcome

OpenStudy (sshayer):

write it like \[\sum_{n=1}^{16}\left\{ -9+\left( n-1 \right)6 \right\}\] \[=\sum_{n=1}^{16}\left( 6n-15 \right)\]

OpenStudy (studyloverhojea):

-9 - 3 + 3 + 9 + ... + 81 Write it following the order of PEMDAS.